### v1

# Navigating PESQ: Up-to-Date Versions and Open Implementations

Abstract
Perceptual Evaluation of Speech Quality (PESQ) is an objective quality measure that remains widely used despite its withdrawal by the International Telecommunication Union (ITU). PESQ has evolved over two decades, with multiple versions and publicly available implementations emerging during this time. The numerous versions and their updates can be overwhelming, especially for new PESQ users. This work provides practical guidance on the different versions and implementations of PESQ. We show that differences can be significant, especially between PESQ versions. We stress the importance of specifying the exact version and implementation that is used to compute PESQ, and possibly to detail how multi-channel signals are handled. These practices would facilitate the interpretation of results and allow comparisons of PESQ scores between different studies. We also provide a repository that implements the latest corrections to PESQ, i.e., Corrigendum 2, which is not implemented by any other openly available distribution: https://github.com/audiolabs/PESQ.
1 Introduction
1.1 PESQ history, importance, and caveats
PESQ [1, 2, 3, 4, 5] was developed to predict the mean opinion score (MOS) of a transmitted speech signal by comparing it with a clean reference. PESQ has a long history as an International Telecommunications Union – Telecommunications Sector (ITU-T) recommendation P.862, with its first version published in 2001 and the last document dated 2018, as summarized in Table 1. In 2018, PESQ was withdrawn by the ITU to be replaced by Perceptual Objective Listening Quality Analysis (POLQA) [6, 7]. Despite its withdrawal, PESQ remains a widely used metric for researchers and practitioners. This is evident by conducting a Google Scholar search for the term PESQ, which returns over 4,600 search results for the year 2024, while only a little more than 100 results are found for the term POLQA. More recently, methods have been proposed to predict PESQ scores non-intrusively (i.e., without a clean reference) using a neural network [8, 9]. The popularity of PESQ is due to several reasons; chief among them is the abundance of publicly available implementations and the robustness that PESQ shows across different application domains [10]. Still, given the evolution of PESQ and its versions, the availability of several implementations can lead to confusion and may render PESQ values reported in different works incomparable.
1.2 Contributions and structure of this work
This work examines PESQ implementations and their differences within the larger context of ITU recommendations. The PESQ versions and implementations considered in this work are summarized in Table 2. The remainder of this work is structured as follows. Sec. 1.3 details the test audio signals, Sec. 2 describes the different versions of PESQ. Sec. 3 discusses the different publicly available implementations. The application to multi-channel signals is discussed in Sec. 4. Finally, additional remarks (Sec. 5) and conclusions (Sec. 6) complete this work.
1.3 Test audio signals
As test audio signals in this paper, we use the Open Dataset for Audio Quality (ODAQ) [11]. It is a dataset of stereo audio signals representing a heterogeneous range of processing types and quality levels. In the following experiments both items with and without speech are considered, with the exception of the correlation analysis with subjective scores in Sec. 4, where only signals with speech are selected. For the experiments in Sec. 3, stereo signals are passively downmixed (i.e., by averaging the two channels) before computing PESQ. In all cases, the signals are downsampled to $16$ kHz using librosa [12] with its default parameters, i.e., soxr_hq.
2 PESQ Versions
2.1 Narrowband raw PESQ (P.862)
PESQ was designed to evaluate the quality of speech signals transmitted over telecommunication networks. The method includes a pre-processing step that mimics a telephone handset. Measures for audible disturbances are computed depending on the loudness of the signals and combined into raw PESQ scores according to P.862 [2], i.e., the first version of PESQ.
P.862 (02/2001)
Amendment 1 (03/2003)
P.862 [2] Amendment 2 (11/2005)
Corrigendum 1 (10/2007)
Corrigendum 2 (03/2018) [5]
P.862.1 [3] P.862.1 (11/2003)
P.862.2 (11/2005)
P.862.2 [4] P.862.2 (11/2007)
Corrigendum 1 (10/2017)
Table 1: List of ITU publications defining the different versions of PESQ. In addition to the listed publications, P.862.3 (2005, 2007, and 2011) provides remarks on how to apply P.862, P.862.1, and P.862.2. The documents can be found at https://www.itu.int/rec/T-REC-P/en.
Implementation PESQ version Narrowband raw PESQ (P.862) 8 kHz / 16 kHz Narrowband MOS-LQO (P.862.1) 8 kHz / 16 kHz Wideband MOS-LQO (P.862.2) 16 kHz without Corrigendum 2 Wideband MOS-LQO (P.862.2) 16 kHz with Corrigendum 2 Accepts stereo inputs ITU reference code (ANSI C) (RMSE = 0.58) (RMSE = 0.57) (Reference) (RMSE = 0.52) (Intended?) ludlows/PESQ (Python) (RMSE = 0.03) audiolabs/torch-pesq (PyTorch) (RMSE = 0.14) vBaiCai/python-pesq (Python) Loizou (MATLAB) audiolabs/PESQ (Python)
Table 2: Overview of the main PESQ versions and their publicly available implementations. The PESQ versions are listed in columns, with software implementations marked by a check (✔). Our audiolabs/PESQ provides wideband MOS-LQO (P.862.2) with Corrigendum 2, which is not implemented in the other openly available distributions. Differences between versions can be substantial, as reported by the RMSE with respect to the ITU reference implementation P.862.2 without Corrigendum 2.
2.2 Narrowband MOS-LQO (P.862.1)
A second version was recommended in P.862.1 [3], in which raw PESQ scores are mapped to MOS-LQO (Mean Opinion Score - Listening Quality Objective [13]), as shown in Fig. 1. Both P.862 and P.862.1 support sampling frequencies of 8 and 16 kHz and assume narrow-band telephone-like listening conditions, i.e., with a strong attenuation below 300 Hz and above 3100 Hz.
Figure 1: Mapping from narrowband raw PESQ (P.862) to narrowband MOS-LQO (P.862.1).
2.3 Wideband extension (P.862.2)
The third version of PESQ is P.862.2 [4]. This considers a larger frequency range (50-7000 Hz, named wideband) and is available only for signals with a sampling frequency of 16 kHz. This version only outputs MOS-LQO scores and no raw PESQ scores. Compared to P.862.1, P.862.2 introduces new input filters and a new output mapping to MOS-LQO. Fig. 2 compares narrowband raw PESQ (P.862) to wideband P.862.2 scores, using a sampling frequency of 16 kHz. Differences can be as large as 1.86 with RMSE $=0.58$ on the considered data.
2.4 P.862 Corrigendum 2
More than a decade after the first edition of P.862.2, the filter coefficients for wideband assessment were acknowledged to be incorrect in P.862 Corrigendum 2 [5]. The corrigendum describes a systematic under-prediction of subjective scores. The under-prediction of P.862.2 is estimated to be $0.8$ on average, and the corrected filter coefficients are provided. Fig. 3 shows the differences between P.862.2 with and without Corrigendum 2. The differences in absolute terms are significant with an RMSE $=0.52$ and a maximum difference of $1.30$ for the audio signals considered.
Despite the acknowledged errors, P.862.2 without Corrigendum 2 seems to have established itself as the de facto standard in the audio signal processing community. This could be explained by the lack of an openly available implementation that directly implements Corrigendum 2. This work fills this gap by introducing a publicly available implementation, audioalabs/PESQ, as explained in the following.
Figure 2: Raw PESQ scores (P.862) compared with wideband PESQ (P.862.2 without Corrigendum 2). Reporting Pearson’s correlation $\rho$, RMSE, mean and max difference.
Figure 3: Wideband PESQ (P.862.2) with and without Corrigendum 2. Although the correlation is strong, differences can be as big as 1.30, and RMSE is 0.52 on the considered audio signals.
3 PESQ Implementations
This section reviews the main openly available implementations of PESQ.
3.1 ITU reference implementation (ANSI-C)
The ITU provides a reference ANSI-C implementation [^1] https://www.itu.int/rec/T-REC-P.862-200511-W!Amd2/en that can output all main PESQ versions, with the exception of Corrigendum 2. In fact, Corrigendum 2 specifies these changes in text form, and the actual corrections are left to be implemented by the users of the reference code.
3.2 Python and PyTorch
ludlows/PESQ [^2] https://github.com/ludlows/PESQ wraps around the ITU reference code and provides access to P.862.1 and P.862.2 for Python users. Fig. 4 compares P.862.2 as output by the ITU reference code and by ludlows/PESQ, showing that they are basically identical.
audiolabs/PESQ [^3] https://github.com/audiolabs/PESQ is the repository we created by forking from ludlows/PESQ and adding the corrections detailed in Corrigendum 2. It is released as PyPI package pesqc2. This provides the possibility of easily and reliably computing P.862.2 with Corrigendum 2. The introduction of Corrigendum 2 results in large differences compared to ludlows/PESQ as shown in Fig. 3.
audiolabs/torch-pesq [^4] https://github.com/audiolabs/torch-pesq provides a torch implementation of P.862.2 without Corrigendum 2. As depicted in Fig. 5 and as acknowledged by the authors, some items can deviate significantly from the ITU reference implementation. The average error is rather small with RMSE $=0.14$ on the data considered.
vBaiCai/python-pesq [^5] https://github.com/vBaiCai/python-pesq is also known as pypesq and only implements the narrowband raw PESQ (P.862). Therefore, this package is not suitable for users interested in the more popular wideband PESQ version (P.862.2).
3.3 MATLAB
A MATLAB implementation is available from the CD accompanying Loizou’s textbook [14] (here we consider the 1st edition), and also online as obfuscated, execute-only code [^6] https://ecs.utdallas.edu/loizou/speech/software.htm. Both provide narrowband raw PESQ (P.862) only. Therefore, these packages are also not suitable if wideband PESQ (P.862.2) is of interest.
4 Multi-Channel
4.1 Reference implementation
The results shown so far are obtained considering mono signals. Indeed, PESQ is developed for mono inputs, and all considered publicly available implementations accept only mono inputs (or batches thereof), with one important exception: The ITU reference code accepts multi-channel inputs. It is unclear if this is intentional. In the reference code, the reading routines load all input file samples. In practice, when a multi-channel audio file is given as input, a mono signal is created, in which channels are interleaved. As a consequence, the reference implementation processes the interleaved multi-channel signals as a long single-channel signal.
Figure 4: ludlows/PESQ compared with the ITU reference implementation. Differences are neglectable.
Figure 5: audiolabs/torch-pesq compared with the ITU reference implementation. Some items deviate significantly from the reference implementation.
4.2 Comparing strategies for applying PESQ to stereo signals
In the following, the processing of stereo signals by the ITU reference code is referred to as stereo. More intuitive strategies for applying PESQ to stereo signals include:
  1. 1.
    Mono-downmixing the stereo audio signal before computing PESQ (mono dmx);
  2. 2.
    Evaluating PESQ for the individual channels and then averaging the resulting scores (avg scores).
mono dmx and avg scores are compared in Fig. 6, while mono dmx is compared with stereo in Fig. 7. To better understand the performance of the different strategies for applying PESQ to stereo signals, we analyze the correlation with subjective scores of perceived audio quality. In all cases, we consider the ITU reference implementation of P.862.2, with and without adding the fixes of Corrigendum 2 (c2). As the ground truth, we use the scores collected from two listening tests involving stereo signals:
  1. 1.
    The Dialogue Enhancement (DE) subset of ODAQ, containing speech signals panned to the center and mixed with movie-like music and effects. Speech enhancement methods are applied to boost the level of speech. The considered enhancement methods create significantly different levels of quality, which are evaluated by expert listeners. The remaining non-speech items from ODAQ are not considered, nor are the extra scores from [15].
  2. 2.
    As a second dataset, we consider the results from the listening test described in [16] and named Dialogue Separation (DS). This work compares different speech enhancement methods to improve the level of speech in studio-like recordings.
The correlation coefficients between the different stereo realizations of PESQ and the subjective speech quality scores are provided in Table 3. Both Pearson’s linear correlation and Kendall’s rank correlation are considered. Kendall’s coefficients are mapped as explained in [10, 17] to make their values more comparable to Pearson’s coefficients.
PESQ provides sensible scores in all cases, with correlation coefficients $\geq 0.73$. This prevents drawing a definitive conclusion on the overall best strategy. The relative good performance of PESQ on the stereo signals of ODAQ DE was also observed in [15], where a comparison to other state-of-the-art objective metrics showed PESQ among the best-performing ones. It is worth noting that P.862.2 results in similar correlation with and without Corrigendum 2. As also observed in Fig. 3, Corrigendum 2 introduces a significant difference in absolute terms, rather than in correlation.
Figure 6: Evaluating each channel independently with the ITU reference implementation and then averaging the output PESQ scores compared with passively mono-downmixed inputs.
Figure 7: Stereo signals input to the ITU reference implementation compared with passively mono-downmixed inputs.
5 PESQ for Training
Some works use PESQ as part of the training strategy or validation metrics for Deep Neural Network (DNN)-based speech enhancement (SE), e.g., [18, 19]. In this case, extra care should be taken when using PESQ for the performance evaluation of the trained DNN. For example, in [20], a listening test is reported, which compares SE methods, including a generative method optimized on PESQ. When we compare the subjective scores of this listening test and PESQ P.862.2 outputs, we observe much weaker correlation coefficients than the ones reported so far. Pearson’s correlation is $0.28$, and mapped Kendall’s correlation is $0.31$. This can be explained by the fact that PESQ was overfitted (or fooled). Improvements in PESQ scores did not correspond to improvements in perceptual quality. The problem of DNN-based SE systems that provide high PESQ scores but low perceptual quality is also reported in [21]. In [22], a DNN called PESQNet is proposed to non-intrusively estimate PESQ scores, serving as a differentiable perceptual loss function for optimizing a SE model. To avoid overfitting, the method alternates training between the SE model and PESQNet, effectively guiding the SE model to achieve improved perceptual quality measured by higher PESQ scores.
ODAQ DE [11] DS [16]
ITU ref. P.862.2 (stereo) $73$ $66$ $83$ $85$
ITU ref. P.862.2 (mono dmx) $80$ $73$ $83$ $87$
ITU ref. P.862.2 (avg scores) $76$ $68$ $85$ $90$
ITU ref. P.862.2 c2 (stereo) $74$ $75$ $85$ $88$
ITU ref. P.862.2 c2 (mono dmx) $79$ $77$ $81$ $88$
ITU ref. P.862.2 c2 (avg scores) $75$ $71$ $83$ $90$
Table 3: Pearson’s and mapped Kendall’s correlation between subjective scores from listening tests on stereo speech enhancement and different strategies for applying PESQ to stereo signals.
6 Conclusion
While PESQ is a widely used speech quality metric, the term PESQ encompasses different proposals and versions. These include: narrowband raw PESQ score (P.862), narrowband MOS-LQO (P.862.1), and wideband MOS-LQO (P.862.2), with and without Corrigendum 2. This corrigendum introduced significant differences, but was not included in any of the numerous publicly available implementations of PESQ. Probably due to this lack, the community seemed to ignore the corrigendum in practice, and P.862.2 without Corrigendum 2 was used most frequently. Our audiolabs/PESQ fills this gap and provides P.862.2 with Corrigendum 2, i.e., the latest version of PESQ.
We also provided practical guidance on the different versions and implementations. We observed that the differences between PESQ versions are generally much more significant than those between different implementations of the same version. A challenge emerges when dealing with multi-channel signals. The reference PESQ implementation by the ITU processes stereo signals and outputs plausible scores. Other more intuitive extensions for stereo processing are discussed.
Finally, it is recommended to specify the PESQ version and implementation when reporting PESQ scores, as this ensures clarity and reproducibility across studies.